satellite73 (satellite73):

what the monkey are "parent functions" and what math teacher made this stuff up??

Vocaloid (vocaloid):

like, say you have something like f(x) = 5(x-4)^2 + 2

Vocaloid (vocaloid):

the "parent function" is the function with all the translations/dilations/etc. removed

Vocaloid (vocaloid):

so the parent function of f(x) is just x^2

OpenStudy (loser66):

Yes, I was taught exactly this.

satellite73 (satellite73):

@Loser66 for real?

OpenStudy (loser66):

Yes, for many functions like e^..., ln.... and trigs

jimthompson5910 (jim_thompson5910):

parent functions are functions like y = x^2 or y = sin(x) The basic form of the function the "child" function is manipulations of the parent function parent function: y = x^2 child function: y = (x-3)^2 .... shifted parent function 3 units to the right another child function: y = x^2+5 ... shift parent function up 5 units